OpenStudy (anonymous):

which of the following are equal to the expression below . 2* square root 36 a. sqrt 144 b.72 c.sqrt 72 d.144 e.12 f.36 ?? help plz some one

OpenStudy (anonymous):

a and e

OpenStudy (hoblos):

2sqrt36 = sqrt(4*36)=sqrt(144) = 12